Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎

day12


Crew 100A Log Book for February 23, 2011

Photos of the day





Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA:

EVA #14 (09:30-10:00),

EVA#15 (14:35-15:04),

EVA#16 (15:33-15:57),

EVA#17 (17:30-18:00),

EVA#18 (17:00-18:20)

Brief Narrative of Field Mission Results:

During the morning we finished the habitability experiments and conducted an EVA to collect extra samples on P2 (N38.40746º W110.79280º) and P3  (N38.40737º W110.79261º) by foot for our cooperation with the Vrije Universiteit of Amsterdam. The samples consisted of small rocks and soil taken at different positions and at different depths and will be sent once we get back to Europe along with the ones for the ESTEC and the SALM group in Reunion.

The afternoon was reserved for another 4 EVAs, 2 EVAs as part of the experiment to monitor the effect of communication-induced stress on an EVA (both by foot), another EVA to get more details on the P3 position (by foot) and finally another EVA by ATV to collect samples at the location C4 (N38.40545º W110.78573º, AKA Dino site near the Habitat - near a pale green vertical fossilized tree trunk) and location P6 (N38.42638º W110.78342º).

Hopefully tonight there will be some time for star observation in the small mountain near the Habitat.

Engineering/Hab Maintenance: GreenHab system is working. water has been manually pumped but needs some time to get through the straining tank to be fully operational.

Plans for Tomorrow:  Continue with more EVAs to test the influence of stress induced by communication on the EVAs, explore the possibility of using the range extender with the parrot.

Reports scheduled for today: Engineering, Commander, Journalist, Science, EVA #14, EVA#15, EVA#16, EVA#17 and EVA#18

Inventory: N/A

Support Requested: We only have about 1.5 containers of gas left for the ATVs. Consider bringing some gas during crew change over.

Miscellaneous: N/A



Commander's Report

Ivo Ferreira, Crew 100A, Commander

Digging Mars

After finishing the last habitability experiments we now have a lot of time to perform EVAs. We are becoming "EVA experts" here on Mars, today we managed to complete 5 of them: two to continue the communication-induced stress, and 3 more to collect extra samples for the ESTEC and the Vrije Universiteit of Amsterdam.

The engineers back on Earth did a great job designing our Hab but what we really like is to be outside to be able to do some science and at the same time enjoy the magnificent landscape of this planet.

During the afternoon alone, we did the two EVAs with simulated emergencies and gathered about 2 kg of samples at different locations (don't worry there is still a lot of sand, rock and dirt here on Mars), and for this we mostly used the ATVs.

They are also great pieces of engineering. They allow us to go faster, further and spend much less energy that would otherwise turn our experience into an Extremely Violent Activity. Some would say they look like the quads on Earth but they are much more fun. As we get more used to them we find that every trip that is more than a few hundred meters away is a good excuse to use them...

The afternoon ended with Audrey testing the range of Max (which was already missing us) with and without the the range extender. We have to get to the bottom of the problem and understand why it does not seem  to work when the camera is broadcasting. We understand that Max is a lady and she wants her privacy.

Tonight we will enjoy some curry rice and hopefully stare at the sky, the weather is clean so we can probably see Earth from here tonight.


Engineering Report

Kévin Theatre Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.75

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.50

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 6 Gallons


Notes/Comments: We have planned to have EVAs with ATVs in the last days. We might be short of gasoline.


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 10

· Trailer Potable Water Tank Level (inches from bottom): Full

· Hab Potable Water Tank Level (inches from bottom): 9

· Potable Water Meter Reading: 424600


Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: Green water

· Condition of plants in Tank 2: Ggreen things

· Condition of plants in Tank 3: Green things

· Greenhab notes/ comments: The level of water of the 3 tanks has reduced a little. The process is slowly getting into place.


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): no

· Oil Checked (yes/no): No

· Tire Status: checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car: Just to supply in electricity the relay for the Rover during our first attempt to do mission SALM.


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: Yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status: filled

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: No

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: yes

· Oil Checked: yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 68

· Thermostat setting downstairs (Farenheit): 60

· Actual temperature downstairs (Farenheit): 58


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:


Science Report

Ayako Ono Reporting

1. Habitability Experiment: Today is the final day for the experiment to be performed with stimuli. This morning, Ivo, Kévin, and Ayako discussed about color wheel, first. Currently, we cannot use printer with our laptops because of the version, and I decided to use a package of crackers and color cards as freshwater checker, instead of color wheel. I took a picture of materials as color gradation strips together with them and white paper. Please see the attached file. I will try to take these photos under ideal white light, sometime soon, so that I will be able to adjust the color of color gradations, and it will work instead of color wheel. After the discussion, Audrey performed with “Copy and Mirror” as neutral stimulation, which is not so interesting task as copy, and Kévin performed with color strips. Also, habitability questionnaire, as digital file, were offered to all crew members to be filled in. End questionnaire and debriefing are scheduled on 24th Feb.

2. Also in the morning, Audrey and Romain worked on Audrey’s related tests about human factor in tele-operation. She already had a track with a rover, and she tested with the rover and drone, if it improves the situation, as assistant to get information, or not.


Thanks a lot for the suggestions to adjust colors for habitability experiments. They were very helpful and I really appreciated it. Now, you don’t have to send color wheel.


EVA 14 Report

Romain Benchenafi Reporting

EVA # 14, 09:30 - 10:00

Crew members : BENCHENAFI Romain

Site location : P3 UTM 12S N38.40737 W110.79261 

Transit mode : Foot

Objective : Collect samples for ESTEC and SALM

Result : Samples retrieved

Lesson learned : It is easier to collect samples when there is no wind.


EVA 15 Report

Kévin Theatre Reporting

EVA #15, 14:35-15:05 

Crew Members: Ivo Ferreira, Ono Ayako, Kévin Théatre  Site Location: 0518210 4250507 ; 0518272 4250621 ; 0518153 4250453 

Transit Mode: Foot 

Objectives: Mission 5 (testing the influence of stress related to communication on an EVA). 

Results: Managed to comply with the 2 critical objectives. We didn’t manage to find the emergency one. 

Lessons Learned: Having a lot of fog is something very irritating which prevent you from finding some points.


EVA 16 Report

Audrey Lan San Reporting

EVA #16, 23th February 2011, 15:40 - 16:10  

Crew members : Romain Benchenafi and Audrey Lan San

Site location :   

Point India : N 38. W 110.7                         

Point Juliet : N W 110.                         

Emergency point :  

Transit mode : Foot 

Objective : Determine factors of stress during an EVA.  Result : We succeeded in both points India and Juliet and found the point of the emergency.  

Lesson learned : Spare areas when the location is reached to improve the efficiency of research.  


EVA 17 Report

Romain Benchenafi Reporting

EVA # 17, 17:30 - 18:00

Crew members : BENCHENAFI Romain, LAN SAN Audrey

Site location :

P3 UTM 12S N38.40737 W110.79261

P2 UTM 12S  N38.40746 W110.79280  

Transit mode : Foot

Objective : Collect samples for ESTEC and SALM

Result : Samples retrieved

Lesson learned : Working in team is profitable as we can share the tasks


EVA 18 Report

Kévin Theatre Reporting

EVA #18, 17:00-18:20

Crew Members: Ivo Ferreira, Kévin Théatre 

Site Location:

C4 (N38.40737° W110,79261°)

P6 (N38,42638 W110,78342) 

Transit Mode: ATVs 

Objectives: Mission SALM (extraction of samples from C4 et P6). 

Results: Managed to extract the samples for the mission

Lessons Learned: None


Chef's Report

Audrey Lan San Reporting

1a) Today was -

A cooking day  

1b) Was today a special day celebrated at the main meal with special food or activity?

No 

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Brown rice with curry, ginger and garlic. Tuna with mayonnaise. Creamy potatoes soup with carrots, broccolis and onions.  

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded) 

There was too much rice.  

4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.   

No   

5)  Lessons learned  (if any)  6) Comments/questions for Kim and Jean, the food study investigators.

Nothing.


Journalist Report

Romain Benchenafi Reporting

Two days left now before the shift of crews.

As usually we were very busy, with five EVAs today.

Now we are playing UNO all together to finish this day, wishing Jeffrey could have been with us.